### v1

# Corresponding author(s) Equal contribution

Abstract
DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks. However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs. To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on A-Eval, an application-driven benchmark. By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks. Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models. To further assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications.
Keywords:
DeepSeek application-driven evaluation reasoning enhancements capability boundary
1 Introduction
In recent years, large language models (LLMs) have revolutionized natural language processing (NLP). Models such as OpenAI’s GPT series [11], Alibaba’s Qwen series [14], MetaAI’s Llama series [3], and DeepSeek’s DeepSeek series [1] have not only advanced NLP technologies but also empowered intelligent solutions for real-world applications. Notably, DeepSeek-R1 [2], with its extremely low training cost, has achieved state-of-the-art (SOTA) performance comparable to OpenAI’s models [13]. Meanwhile, its exceptional reasoning capabilities have attracted significant attention.
Comprehensive evaluation is essential for understanding model strengths and limitations, as well as guiding optimization and practical deployment. While DeepSeek-R1 [2] has demonstrated SOTA performance on many existing benchmarks such as MMLU [5], C-Eval [6], SimpleQA [12], LiveCodeBench [7], Math-500 [10], these benchmarks do not fully capture the nuances of real-world applications. Consequently, two critical challenges arise when deploying DeepSeek models in practical scenarios: First, how should users determine the most suitable model scale? While larger models offer better capabilities, they also come with higher deployment and inference costs. Second, are reasoning-enhanced models always better for all tasks? If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by non-enhanced models?
To address the gap, we evaluate DeepSeek models on A-Eval [9], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories. Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks. These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
2 Evaluation Framework
2.1 Evaluated Models
To assess the impact of reasoning enhancements in DeepSeek models for real-world applications, we evaluate 14 models grouped into seven pairs, including the Mixture-of-Experts (MoE) and dense models of various series and scales. Each pair consists of an instruction-tuned model and its corresponding distilled or reasoning-enhanced version, both derived from the same base model. Tab.  1 lists the evaluated models and their descriptions.
Table 1: The Evaluated model list.
Model Description
Qwen2.5-Math-1.5B-Instruct Dense math Model, SFT from Qwen2.5-Math-1.5B
DeepSeek-R1-Distill-Qwen-1.5B Dense model distilled from Qwen2.5-Math-1.5B
Qwen2.5-Math-7B-Instruct Dense math Model, SFT from Qwen2.5-Math-7B
DeepSeek-R1-Distill-Qwen-7B Dense model Distilled from Qwen2.5-Math-7B
Llama-3.1-8B-Instruct Dense model SFT from Llama-3.1-8B-Base
DeepSeek-R1-Distill-Llama-8B Dense model Distilled from Llama-3.1-8B-Base
Qwen2.5-14B-Instruct Dense model SFT from Qwen2.5-14B-Base
DeepSeek-R1-Distill-Qwen-14B Dense model Distilled from Qwen2.5-14B-Base
Qwen2.5-32B-Instruct Dense model SFT from Qwen2.5-32B-Base
DeepSeek-R1-Distill-Qwen-32B Dense model Distilled from Qwen2.5-32B-Base
Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-7B-Base
DeepSeek-R1-Distill-Llama-70B Dense model Distilled from Llama-3.3-70B-Instruct
DeepSeek-V3 (671B|37B) MoE model SFT from DeepSeek-V3-Base
DeepSeek-R1 (671B|37B) MoE model cold-start + RL from DeepSeek-V3-Base
2.2 Dataset
The A-Eval benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across five major categories and 27 subcategories. We use A-Eval’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications. This evaluation helps assess how DeepSeek-R1’s reasoning enhancements improve model capabilities in different practical scenarios.
2.3 Evaluation Process
We follow A-Eval’s zero-shot automatic evaluation process, which includes:
Inference: Feed each question ${{Q_{i}}}$ into the evaluated model to generate a prediction ${{P_{i}}}$.
Triplet Preparation: Construct the triplet $\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, where ${{A_{i}}}$ is the ground-truth answer for ${{Q_{i}}}$.
Scoring: Combine the prompt and triplet$\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, input them into the scoring model, and obtain the scoring output ${{S_{i}}}$ between 0 and 100. We use A-Eval’s scoring prompt and a greater LLM, Qwen2.5-72B-Instruct as the scoring model.
3 Results and Discussion
3.1 Overall Performance
Fig. 1 (a) presents the overall average scores of each model across all data.
Findings Consistent with Common Sense:
(1) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts.
(2) Overall, DeepSeek-V3 and DeepSeek-R1 demonstrate superior performance compared to other model families.
(3) Within the same series, models follow the scaling law [4, 8] regardless of whether they have undergone reasoning enhancement.
(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 178.74% and 54.36%, respectively. This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
Figure 1: Average score of DeepSeek series models on A-Eval. (a) The overall average score of models across all data. (b) to (f). The average scores of models on each task. The "Instruct Model" refers to the original models without reasoning enhancement, and the "Reasoning Enhancement Model" refers to models that have been enhanced using DeepSeek-R1’s reasoning data.
Findings Contrary to Common Sense:
(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B.
(6) Qwen-2.5-14B exhibits performance degradation after distillation.
3.2 Performance Comparison By Task
Fig. 1(b) to Fig. 1 (f) compare the scores of the models on the five major task categories.
Findings Consistent with Common Sense:
(7) The two mathematical models achieve comparable performance in Logical Reasoning but under-perform in other tasks.
(8) For Logical Reasoning tasks, all models show improvement after distillation, with the increase being greater than for other tasks.
Findings Contrary to Common Sense:
(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3). For Task Planning, apart from the two mathematical models, other models either remain unchanged or see a decline in performance after distillation.
(10) For Logical Reasoning tasks, original Llama-3.3-70B underperforms the smaller model Qwen2.5-14B. After distillation, DeepSeek-R1-Distilled-Llama-3.3-70B surpasses DeepSeek-R1-Distilled-Qwen2.5-32B.
(11) For Information Extraction tasks, Qwen2.5-32B consistently outperform the larger model Llama-3.3-70B both before and after distillation.
Figure 2: Average scores of the DeepSeek series models on the 27 subcategories.
3.3 Performance Comparison By Subtask
In more detail, Fig. 2 provides the scores of the models on the 27 subcategories.
After analyzing these results, here are some interesting findings:
(12) DeepSeek models dominate 23/27 subtasks, except for Short Text Classification, Named Entity Recognition, and Common Sense Question Answering.
(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-11.3%), Part-of-Speech Tagging (-13.4%), and Open-source Question Answering (-17.05%).
(14) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 31.45%.
(15) For arithmetic operations tasks, distillation provides benefits to all models except for Qwen2.5-Math-7B (85.26 to 84).
3.4 Performance Comparison By Model
To more clearly compare the performance of each group of models before and after reasoning enhancement on the five major tasks, we present the evaluation scores by seven model groups in Fig. 3.
Figure 3: Performance of each model group on five major tasks.
3.5 Model Selection Guidance for Users
To assist users in selecting the most suitable DeepSeek model based on their specific application requirements, we present the evaluation results using line charts, following the A-Eval methodology. These visualizations provide a clear and intuitive comparison of model performance across different tasks and subcategories.
Figure 4: Line chart for model selection on five major tasks.
Figure 4 illustrates the scores of the 14 evaluated models on each major task, as well as their average scores across all tasks. This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
In Fig. 5, we further break down the performance of the 14 models by showing their average scores within each of the five major task categories and their corresponding subcategories. These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
Figure 5: Line chart for model selection on five major tasks and corresponding subtasks.
Using these capability boundary quantization curves, users can make informed decisions when selecting the most suitable DeepSeek model for their specific tasks and performance expectations.
Table 2: Performance tier classification of DeepSeek models across major task categories. Models are categorized into tiers (A+: >85, A: 80-85, B: 70-80, C: 60-70, D: <60) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
Model Model Scale TU IE TG LR TP
Qwen2.5-Math-1.5B-Instruct 1.5B D D D D D
DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D C D
Qwen2.5-Math-7B-Instruct 7B D D D C D
DeepSeek-R1-Distill-Qwen-7B 7B C D C B B
Llama-3.1-8B-Instruct 8B C C C D B
DeepSeek-R1-Distill-Llama-8B 8B C C B C B
Qwen2.5-14B-Instruct 14B B B A B A+
DeepSeek-R1-Distill-Qwen-14B 14B B B B A A
Qwen2.5-32B-Instruct 32B B B A B A
DeepSeek-R1-Distill-Qwen-32B 32B B B A B A+
Llama-3.3-70B-Instruct 70B B B B B A
DeepSeek-R1-Distill-Llama-70B 70B B B B A A
DeepSeek-V3 671B|37B A B A+ A A+
DeepSeek-R1 671B|37B A A A A+ A
To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories. As shown in Tab. 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 60). This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
Based on Tab. 2, we provide specific examples of model selection to illustrate its practical utility:
If users require A-level performance, the following models are recommended for each task category (highlighted in red in Tab. 2):
Text Understanding: DeepSeek-V3 (A)
Information Extraction: DeepSeek-R1 (A)
Text Generation: Qwen2.5-14B-Instruct (A)
Logical Reasoning: DeepSeek-R1-Distill-Qwen-14B (A)
Task Planning: Qwen2.5-14B-Instruct (A+)
For users with more flexible requirements, such as B-level performance, the following models offer a cost-effective balance (highlighted in green in Tab. 2):
Text Understanding: Qwen2.5-14B-Instruct (B)
Information Extraction: Qwen2.5-14B-Instruct (B)
Text Generation: DeepSeek-R1-Distill-Llama-8B (B)
Logical Reasoning: DeepSeek-R1-Distill-Qwen-7B (B)
Task Planning: DeepSeek-R1-Distill-Qwen-7B (B)
These examples demonstrate how the tiered classification can guide users in selecting models that meet their specific performance and cost requirements, ensuring optimal deployment in real-world applications.
4 Conclusion
To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the A-Eval benchmark. Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks. Finally, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts. Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements. In future work, we plan to expand the tasks and datasets of A-Eval to offer even more comprehensive insights for model selection and application.

### v2

@@
-Corresponding author(s) Equal contribution
+Corresponding author(s)
 Abstract
@@
 However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs.
-To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on A-Eval, an application-driven benchmark.
+To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on the improved version A-Eval (A-Eval-2.0), an application-driven benchmark.
 By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks.
@@
 Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications.
+It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results.
+We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results.
+Please refer to the latest version of the paper for the most recent results and conclusions.
 Keywords:
@@
 If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by non-enhanced models?
-To address the gap, we evaluate DeepSeek models on A-Eval [9], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories.
+To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [9], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories.
+Then, we evaluate DeepSeek models on A-Eval-2.0.
 Through this evaluation, we extend existing evaluations and provide actionable insights into how reasoning enhancements influence model performance across various practical tasks.
 These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
+It is important to emphasize that the selection of test samples and the design of evaluation criteria will inevitably introduce certain biases into the evaluation results.
+To address this, we will continue to optimize this evaluation work to enhance its comprehensiveness and reliability as much as possible.
 2 Evaluation Framework
@@
 DeepSeek-R1-Distill-Qwen-32B Dense model Distilled from Qwen2.5-32B-Base
-Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-7B-Base
+Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-70B-Base
 DeepSeek-R1-Distill-Llama-70B Dense model Distilled from Llama-3.3-70B-Instruct
@@
 The A-Eval benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across five major categories and 27 subcategories.
-We use A-Eval’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications.
+To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation.
+We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications.
 This evaluation helps assess how DeepSeek-R1’s reasoning enhancements improve model capabilities in different practical scenarios.
 2.3 Evaluation Process
-We follow A-Eval’s zero-shot automatic evaluation process, which includes:
+We follow the automatic and manual scoring methods of A-Eval, and the automatic scoring process includes:
 Inference: Feed each question ${{Q_{i}}}$ into the evaluated model to generate a prediction ${{P_{i}}}$.
@@
 (3) Within the same series, models follow the scaling law [4, 8] regardless of whether they have undergone reasoning enhancement.
-(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 178.74% and 54.36%, respectively.
+(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 169% and 53.6%, respectively.
 This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
-Figure 1: Average score of DeepSeek series models on A-Eval.
+Figure 1: Average score of DeepSeek series models on A-Eval-2.0.
 (a) The overall average score of models across all data.
@@
 Findings Contrary to Common Sense:
-(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B.
+(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B. This may be attributed to the fact that our evaluation dataset includes many Chinese data.
 (6) Qwen-2.5-14B exhibits performance degradation after distillation.
@@
 Findings Consistent with Common Sense:
-(7) The two mathematical models achieve comparable performance in Logical Reasoning but under-perform in other tasks.
+(7) The two mathematical models achieve comparable performance in Logical Reasoning and Task Planning but under-perform in other tasks.
 (8) For Logical Reasoning tasks, all models show improvement after distillation, with the increase being greater than for other tasks.
 Findings Contrary to Common Sense:
-(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3).
-For Task Planning, apart from the two mathematical models, other models either remain unchanged or see a decline in performance after distillation.
+(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3), and Task Planning (Llama-3.1-8B, Llama-3.3-70B).
 (10) For Logical Reasoning tasks, original Llama-3.3-70B underperforms the smaller model Qwen2.5-14B.
@@
 (12) DeepSeek models dominate 23/27 subtasks, except for Short Text Classification, Named Entity Recognition, and Common Sense Question Answering.
-(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-11.3%), Part-of-Speech Tagging (-13.4%), and Open-source Question Answering (-17.05%).
+(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-9.37), Part-of-Speech Tagging (-10.62), and Open-source Question Answering (-12.38%).
 (14) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 31.45%.
-(15) For arithmetic operations tasks, distillation provides benefits to all models except for Qwen2.5-Math-7B (85.26 to 84).
 3.4 Performance Comparison By Model
@@
 Table 2: Performance tier classification of DeepSeek models across major task categories.
-Models are categorized into tiers (A+: >85, A: 80-85, B: 70-80, C: 60-70, D: <60) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
+Models are categorized into tiers (A+: >85, A: 80-85, B: 75-80, C: 65-75, D: <65) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
 Model Model Scale TU IE TG LR TP
 Qwen2.5-Math-1.5B-Instruct 1.5B D D D D D
-DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D C D
-Qwen2.5-Math-7B-Instruct 7B D D D C D
-DeepSeek-R1-Distill-Qwen-7B 7B C D C B B
-Llama-3.1-8B-Instruct 8B C C C D B
-DeepSeek-R1-Distill-Llama-8B 8B C C B C B
-Qwen2.5-14B-Instruct 14B B B A B A+
-DeepSeek-R1-Distill-Qwen-14B 14B B B B A A
-Qwen2.5-32B-Instruct 32B B B A B A
-DeepSeek-R1-Distill-Qwen-32B 32B B B A B A+
-Llama-3.3-70B-Instruct 70B B B B B A
-DeepSeek-R1-Distill-Llama-70B 70B B B B A A
+DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D D C
+Qwen2.5-Math-7B-Instruct 7B D D D D D
+DeepSeek-R1-Distill-Qwen-7B 7B C D D B C
+Llama-3.1-8B-Instruct 8B C D C D B
+DeepSeek-R1-Distill-Llama-8B 8B C D C C C
+Qwen2.5-14B-Instruct 14B B B A B C
+DeepSeek-R1-Distill-Qwen-14B 14B B C B A B
+Qwen2.5-32B-Instruct 32B B B A B C
+DeepSeek-R1-Distill-Qwen-32B 32B B B A B A
+Llama-3.3-70B-Instruct 70B C C C C B
+DeepSeek-R1-Distill-Llama-70B 70B B C B A C
 DeepSeek-V3 671B|37B A B A+ A A+
-DeepSeek-R1 671B|37B A A A A+ A
+DeepSeek-R1 671B|37B A A A A+ A+
 To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories.
 As shown in Tab.
-2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 60).
+2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65).
 This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
 Based on Tab. 2, we provide specific examples of model selection to illustrate its practical utility:
+If users require A+ level performance, the following models are recommended for each task category:
+Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
+Information Extraction: DeepSeek-R1 (A)
+Text Generation: DeepSeek-V3 (A+)
+Logical Reasoning: DeepSeek-R1 (A+)
+Task Planning: DeepSeek-V3 (A+) or DeepSeek-R1 (A+)
 If users require A-level performance, the following models are recommended for each task category (highlighted in red in Tab. 2):
@@
 Logical Reasoning: DeepSeek-R1-Distill-Qwen-14B (A)
-Task Planning: Qwen2.5-14B-Instruct (A+)
+Task Planning: DeepSeek-R1-Distill-Qwen-32B (A)
 For users with more flexible requirements, such as B-level performance, the following models offer a cost-effective balance (highlighted in green in Tab. 2):
@@
 Information Extraction: Qwen2.5-14B-Instruct (B)
-Text Generation: DeepSeek-R1-Distill-Llama-8B (B)
+Text Generation: Qwen2.5-14B-Instruct (A)
 Logical Reasoning: DeepSeek-R1-Distill-Qwen-7B (B)
-Task Planning: DeepSeek-R1-Distill-Qwen-7B (B)
+Task Planning: DeepSeek-R1-Distill-Llama-8B (B)
 These examples demonstrate how the tiered classification can guide users in selecting models that meet their specific performance and cost requirements, ensuring optimal deployment in real-world applications.
+It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria.
+Therefore, our evaluation results may not be the most objective, comprehensive, or accurate.
+As we will continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
 4 Conclusion
-To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the A-Eval benchmark.
+To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the modified A-Eval-2.0 benchmark.
 Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks.
@@
 Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements.
-In future work, we plan to expand the tasks and datasets of A-Eval to offer even more comprehensive insights for model selection and application.
+In the future, we will continue to advance this work by optimizing evaluation benchmarks and promptly updating evaluation results to the community.

### v3

@@
 However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs.
-To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, and DeepSeek-R1-Distill-Llama series on the improved version A-Eval (A-Eval-2.0), an application-driven benchmark.
+To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, and their corresponding 4-bit quantized models on the enhanced A-Eval benchmark, A-Eval-2.0.
 By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks.
@@
 While DeepSeek-R1 [3] has demonstrated SOTA performance on many existing benchmarks such as MMLU [6], C-Eval [7], SimpleQA [15], LiveCodeBench [8], Math-500 [13], these benchmarks do not fully capture the nuances of real-world applications.
-Consequently, two critical challenges arise when deploying DeepSeek models in practical scenarios: First, how should users determine the most suitable model scale?
+Consequently, three critical challenges arise when deploying DeepSeek models in practical scenarios: First, how should users determine the most suitable model scale?
 While larger models offer better capabilities, they also come with higher deployment and inference costs.
@@
 If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by non-enhanced models?
+Third, is the extent of performance loss observed in the quantized model when applied to various real-world tasks within an acceptable range?
 To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [10], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories.
@@
 2.1 Evaluated Models
-To assess the impact of reasoning enhancements in DeepSeek models for real-world applications, we evaluate 14 models grouped into seven pairs, including the Mixture-of-Experts (MoE) and dense models of various series and scales.
-Each pair consists of an instruction-tuned model and its corresponding distilled or reasoning-enhanced version, both derived from the same base model.
-Tab.
-1 lists the evaluated models and their descriptions.
+To assess the impact of reasoning enhancements in DeepSeek models and assist users in identifying the most cost-effective models for practical applications, we conduct a systematic assessment of 21 models organized into seven distinct groups.
+These models encompass both Mixture-of-Experts (MoE) and dense models across various series and scales.
+Each group includes an instruction-tuned model, its corresponding distilled or reasoning-enhanced variant, and the associated quantized model.
+The instruction-tuned models are from HuggingFace  [17, 11], the distilled models are from HuggingFace  [1], and the quantized models are from Ollama  [12].
+Table 1 lists the evaluated models and their descriptions.
 Table 1: The Evaluated model list.
@@
 DeepSeek-R1-Distill-Qwen-1.5B Dense model distilled from Qwen2.5-Math-1.5B
+Distill-Qwen-1.5B-Q4KM The 4-bit quantized version of distilled model
 Qwen2.5-Math-7B-Instruct Dense math Model, SFT from Qwen2.5-Math-7B
 DeepSeek-R1-Distill-Qwen-7B Dense model Distilled from Qwen2.5-Math-7B
+Distill-Qwen-7B-Q4KM The 4-bit quantized version of distilled model
 Llama-3.1-8B-Instruct Dense model SFT from Llama-3.1-8B-Base
 DeepSeek-R1-Distill-Llama-8B Dense model Distilled from Llama-3.1-8B-Base
+Distill-Llama-8B-Q4KM The 4-bit quantized version of distilled model
 Qwen2.5-14B-Instruct Dense model SFT from Qwen2.5-14B-Base
 DeepSeek-R1-Distill-Qwen-14B Dense model Distilled from Qwen2.5-14B-Base
+Distill-Qwen-14B-Q4KM The 4-bit quantized version of distilled model
 Qwen2.5-32B-Instruct Dense model SFT from Qwen2.5-32B-Base
 DeepSeek-R1-Distill-Qwen-32B Dense model Distilled from Qwen2.5-32B-Base
+Distill-Qwen-32B-Q4KM The 4-bit quantized version of distilled model
 Llama-3.3-70B-Instruct Dense model SFT from Llama-3.3-70B-Base
 DeepSeek-R1-Distill-Llama-70B Dense model Distilled from Llama-3.3-70B-Instruct
+Distill-Llama-70B-Q4KM The 4-bit quantized version of distilled model
 DeepSeek-V3 (671B|37B) MoE model SFT from DeepSeek-V3-Base
 DeepSeek-R1 (671B|37B) MoE model cold-start + RL from DeepSeek-V3-Base
+DeepSeek-R1-Q4KM The 4-bit quantized version of DeepSeek-R1
 2.2 Dataset
-The A-Eval benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across five major categories and 27 subcategories.
-To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation.
+The A-Eval [10] benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across three difficulty level, five major categories and 27 subcategories.
+To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the sub-tasks and QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation.
 We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications.
-This evaluation helps assess how DeepSeek-R1’s reasoning enhancements improve model capabilities in different practical scenarios.
+On one hand, the evaluation on A-Eval-2.0 enables researchers to assess how the reasoning enhancements in DeepSeek-R1 improve model performance across various practical scenarios.
+On the other hand, the evaluation results assist users in selecting the most cost-effective models for their specific application needs.
 2.3 Evaluation Process
@@
 3 Results and Discussion
+In this section, we first present the overall average performance of 21 models across all data in A-Eval-2.0 (Section 3.1).
+Next, we analyze the models’ performance from three perspectives: task category (Section 3.2), sub-task category (Section 3.3), and model group (Section 3.4).
+Based on these evaluation results, we provide line charts and tier classification table to assist users in selecting the most cost-effective models for their specific requirements (Section 3.5).
+Finally, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6).
 3.1 Overall Performance
-Fig. 1 (a) presents the overall average scores of each model across all data.
+Figure 1 (a) presents the overall average scores of each model group across all data.
 Findings Consistent with Common Sense:
-(1) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts.
-(2) Overall, DeepSeek-V3 and DeepSeek-R1 demonstrate superior performance compared to other model families.
+(1) Undoubtedly, DeepSeek-V3 and DeepSeek-R1 demonstrate the best overall performance among all evaluated models..
+(2) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts. After 4-bit quantization, the model’s scores decrease by an average of 3.52%.
 (3) Within the same series, models follow the scaling law [5, 9] regardless of whether they have undergone reasoning enhancement.
-(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 169% and 53.6%, respectively.
+(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 178.74% and 54.36%, respectively.
 This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
@@
 The average scores of models on each task.
-The "Instruct Model" refers to the original models without reasoning enhancement, and the "Reasoning Enhancement Model" refers to models that have been enhanced using DeepSeek-R1’s reasoning data.
+The "Instruct Model" refers to the original Instruction-tuned models without reasoning enhancement, the "Reasoning Enhancement Model" refers to reasoning-enhanced models that have been enhanced using DeepSeek-R1’s reasoning data, and the "Quantization Model" refers to 4-bit quantized version of the reasoning enhancement models.
 Findings Contrary to Common Sense:
-(5) Before and after distillation, Qwen2.5-32B outperforms the larger model, Llama-3.3-70B. This may be attributed to the fact that our evaluation dataset includes many Chinese data.
+(5) Both before and after distillation, Qwen2.5-32B consistently outperforms the larger Llama-3.3-70B model.
+The counterintuitive result is likely due to the predominance of Chinese data in our evaluation dataset, and this phenomenon also appears later in the discussion of performance comparison by task (Section 3.2) and by subtask (Section 3.3).
 (6) Qwen-2.5-14B exhibits performance degradation after distillation.
 3.2 Performance Comparison By Task
-Fig. 1(b) to Fig. 1 (f) compare the scores of the models on the five major task categories.
+Figure 1(b) to  Figure 1 (f) compare the scores of the models on the five major task categories.
 Findings Consistent with Common Sense:
@@
 (8) For Logical Reasoning tasks, all models show improvement after distillation, with the increase being greater than for other tasks.
+Additionally, the performance degradation of 4-bit quantization on Logical Reasoning tasks is the most significant, at 7.91%.
+Figure 2: Average scores of the DeepSeek series models on the 27 subcategories.
 Findings Contrary to Common Sense:
@@
 (11) For Information Extraction tasks, Qwen2.5-32B consistently outperform the larger model Llama-3.3-70B both before and after distillation.
-Figure 2: Average scores of the DeepSeek series models on the 27 subcategories.
 3.3 Performance Comparison By Subtask
-In more detail, Fig. 2 provides the scores of the models on the 27 subcategories.
+In more detail, Figure 2 provides the scores of the models on the 27 subcategories.
 After analyzing these results, here are some interesting findings:
 (12) DeepSeek models dominate 23/27 subtasks, except for Short Text Classification, Named Entity Recognition, and Common Sense Question Answering.
-(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-9.37), Part-of-Speech Tagging (-10.62), and Open-source Question Answering (-12.38%).
+(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-11.3%), Part-of-Speech Tagging (-13.4%), and Open-source Question Answering (-17.05%).
 (14) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 31.45%.
 3.4 Performance Comparison By Model
-To more clearly compare the performance of each group of models before and after reasoning enhancement on the five major tasks, we present the evaluation scores by seven model groups in Fig. 3.
+To more clearly compare the performance of each group of models before and after reasoning enhancement on the five major tasks, we present the evaluation scores by seven model groups in Figure 3.
 Figure 3: Performance of each model group on five major tasks.
@@
 Figure 4: Line chart for model selection on five major tasks.
-Figure 4 illustrates the scores of the 14 evaluated models on each major task, as well as their average scores across all tasks.
+Figure 4 illustrates the scores of the 21 evaluated models on each major task, as well as their average scores across all tasks.
 This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
-In Fig.
-5, we further break down the performance of the 14 models by showing their average scores within each of the five major task categories and their corresponding subcategories.
+In Figure 5, we further break down the performance of the 21 models by showing their average scores within each of the five major task categories and their corresponding subcategories.
 These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
@@
 DeepSeek-R1-Distill-Qwen-1.5B 1.5B D D D D C
+Distill-Qwen-1.5B-Q4KM 1.5B D D D D C
 Qwen2.5-Math-7B-Instruct 7B D D D D D
 DeepSeek-R1-Distill-Qwen-7B 7B C D D B C
+Distill-Qwen-7B-Q4KM 7B D D D C C
 Llama-3.1-8B-Instruct 8B C D C D B
 DeepSeek-R1-Distill-Llama-8B 8B C D C C C
+Distill-Llama-8B-Q4KM 8B C D C D C
 Qwen2.5-14B-Instruct 14B B B A B C
 DeepSeek-R1-Distill-Qwen-14B 14B B C B A B
+Distill-Qwen-14B-Q4KM 14B B C A C B
 Qwen2.5-32B-Instruct 32B B B A B C
 DeepSeek-R1-Distill-Qwen-32B 32B B B A B A
+Distill-Qwen-32B-Q4KM 32B B B A+ B B
 Llama-3.3-70B-Instruct 70B C C C C B
 DeepSeek-R1-Distill-Llama-70B 70B B C B A C
+Distill-Llama-70B-Q4KM 70B B C B C C
 DeepSeek-V3 671B|37B A B A+ A A+
 DeepSeek-R1 671B|37B A A A A+ A+
+DeepSeek-R1-Q4KM 671B|37B B C A A+ A+
 To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories.
-As shown in Tab.
-2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65).
+As shown in Table 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65).
 This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
-Based on Tab. 2, we provide specific examples of model selection to illustrate its practical utility:
-If users require A+ level performance, the following models are recommended for each task category:
+Based on Table 2, we provide specific examples of model selection to illustrate its practical utility:
+If users require A+ level performance, the following models are recommended for each task category ((highlighted in blue in Table 2)):
 Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
 Information Extraction: DeepSeek-R1 (A)
-Text Generation: DeepSeek-V3 (A+)
-Logical Reasoning: DeepSeek-R1 (A+)
-Task Planning: DeepSeek-V3 (A+) or DeepSeek-R1 (A+)
-If users require A-level performance, the following models are recommended for each task category (highlighted in red in Tab. 2):
-Text Understanding: DeepSeek-V3 (A)
+Text Generation: Distill-Qwen-32B-Q4KM (A+)
+Logical Reasoning: DeepSeek-R1-Q4KM (A+)
+Task Planning: DeepSeek-R1-Q4KM (A+)
+If users require A level performance, the following models are recommended for each task category (highlighted in red in Table 2):
+Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
 Information Extraction: DeepSeek-R1 (A)
-Text Generation: Qwen2.5-14B-Instruct (A)
+Text Generation: Distill-Qwen-14B-Q4KM (A)
 Logical Reasoning: DeepSeek-R1-Distill-Qwen-14B (A)
 Task Planning: DeepSeek-R1-Distill-Qwen-32B (A)
-For users with more flexible requirements, such as B-level performance, the following models offer a cost-effective balance (highlighted in green in Tab. 2):
-Text Understanding: Qwen2.5-14B-Instruct (B)
+For users with more flexible requirements, such as B level performance, the following models offer a cost-effective balance (highlighted in green in Table 2):
+Text Understanding: Distill-Qwen-14B-Q4KM (B)
 Information Extraction: Qwen2.5-14B-Instruct (B)
-Text Generation: Qwen2.5-14B-Instruct (A)
+Text Generation: Distill-Qwen-14B-Q4KM (A)
 Logical Reasoning: DeepSeek-R1-Distill-Qwen-7B (B)
-Task Planning: DeepSeek-R1-Distill-Llama-8B (B)
+Task Planning: Distill-Qwen-14B-Q4KM (B)
 These examples demonstrate how the tiered classification can guide users in selecting models that meet their specific performance and cost requirements, ensuring optimal deployment in real-world applications.
+3.6 Performance Comparison By Difficulty Level
+In A-Eval and A-Eval-2.0, each QA pair is assigned a difficulty level label: Easy, Medium, or Hard.
+It is evident that models tend to achieve lower scores as the difficulty of the questions increases.
+However, there has been no quantitative analysis of how DeepSeek models perform on questions of varying difficulty levels.
+To address this gap, we separately analyze the evaluation results for each difficulty level and present our findings.
+Figure 6 (a) illustrates the average scores of 21 models across task categories for Easy, Medium, and Hard data.
+Meanwhile, Figure 6 (b) - (f) depict the scores of the 21 models for each task category across the three difficulty levels.
+Figure 6: Average performance of 21 models on Easy, Medium, and Hard data.
+In Figure 7, we present the scores of 21 models on Easy, Medium, and Hard data for each sub-task category.
+Figure 7: Performance of 21 models on Easy, Medium, and Hard data for each task and sub-task.
+In Figure 8, we analyze the percentage score gains brought by reasoning enhancements on Easy, Medium, and Hard data. An interesting observation emerges:
+(14) As the difficulty of the data increases, distillation or reasoning enhancements yield higher performance gains for the models.
+Interestingly, reasoning enhancements can even have a negative impact on simpler problems.
+Figure 8: Performance gain of reasoning enhancement on Easy, Medium, and Hard data.
+(a) Performance gain of each model group across all tasks.
+(b) Performance gain of each task category across all models (excluding math models).
 It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria.
@@
 4 Conclusion
-To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models and their distilled variants on the modified A-Eval-2.0 benchmark.
+To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models, their distilled variants, and 4-bit quantized models on the modified A-Eval-2.0 benchmark.
 Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks.

### v4

@@
 DeepSeek-R1, known for its low training cost and exceptional reasoning capabilities, has achieved state-of-the-art performance on various benchmarks.
-However, detailed evaluations from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs.
-To address this gap, we evaluate the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, and their corresponding 4-bit quantized models on the enhanced A-Eval benchmark, A-Eval-2.0.
-By comparing original instruction-tuned models with their distilled counterparts, we analyze how reasoning enhancements impact performance across diverse practical tasks.
+However, detailed evaluations for DeepSeek Series models from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs.
+To address this gap, we conduct a systematic evaluation of the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, their corresponding 4-bit quantized models, and the reasoning model QwQ-32B using the enhanced A-Eval benchmark, A-Eval-2.0.
+Through a comparative analysis of original instruction-tuned models and their distilled counterparts, we investigate how reasoning enhancements impact performance across diverse practical tasks.
 Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models.
-To further assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts.
-Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models, ensuring optimal performance and resource efficiency in real-world applications.
+To assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications.
+Based on the quantification results, we develop a model selection handbook that clearly illustrates the relation among models, their capabilities and practical applications.
+This handbook enables users to select the most cost-effective models without efforts, ensuring optimal performance and resource efficiency in real-world applications.
 It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results.
 We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results.
-Please refer to the latest version of the paper for the most recent results and conclusions.
+Please refer to the latest version of the paper for the most current results and conclusions.
 Keywords:
-DeepSeek application-driven evaluation reasoning enhancements capability boundary
+DeepSeek application-driven evaluation reasoning enhancements capability boundary model selection handbook
 1 Introduction
@@
 Notably, DeepSeek-R1 [3], with its extremely low training cost, has achieved state-of-the-art (SOTA) performance comparable to OpenAI’s models [16].
-Meanwhile, its exceptional reasoning capabilities have attracted significant attention.
+And its exceptional reasoning capabilities have garnered significant attention within the research community.
+Following the success of DeepSeek-R1, Alibaba introduced QwQ-32B [19], a dense reasoning model that also demonstrates competitive performance.
 Comprehensive evaluation is essential for understanding model strengths and limitations, as well as guiding optimization and practical deployment.
@@
 Second, are reasoning-enhanced models always better for all tasks?
-If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by non-enhanced models?
+If not, which tasks are best suited for reasoning-enhanced models, and which tasks can be handled by original instruction-tuned models?
 Third, is the extent of performance loss observed in the quantized model when applied to various real-world tasks within an acceptable range?
@@
 2.1 Evaluated Models
-To assess the impact of reasoning enhancements in DeepSeek models and assist users in identifying the most cost-effective models for practical applications, we conduct a systematic assessment of 21 models organized into seven distinct groups.
+To assess the impact of reasoning enhancements and quantization in DeepSeek models and assist users in identifying the most cost-effective models for practical applications, we conduct a systematic assessment of 22 models organized into seven distinct groups.
 These models encompass both Mixture-of-Experts (MoE) and dense models across various series and scales.
@@
 DeepSeek-R1-Q4KM The 4-bit quantized version of DeepSeek-R1
+QwQ-32B Dense model with reasoning ability from Qwen2.5-32B-Base
 2.2 Dataset
@@
 3 Results and Discussion
-In this section, we first present the overall average performance of 21 models across all data in A-Eval-2.0 (Section 3.1).
-Next, we analyze the models’ performance from three perspectives: task category (Section 3.2), sub-task category (Section 3.3), and model group (Section 3.4).
-Based on these evaluation results, we provide line charts and tier classification table to assist users in selecting the most cost-effective models for their specific requirements (Section 3.5).
-Finally, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6).
+In this section, we first present the overall average performance of 22 models across all data in A-Eval-2.0 (Section 3.1).
+Subsequently, we analyze the models’ performance from three perspectives: task category (Section 3.2), sub-task category (Section 3.3), and model group (Section 3.4).
+Following the A-Eval methodology, we visualize the comparative evaluation results using line charts (Section 3.5).
+Then, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6).
+Finally, we quantify models’ capability boundary through a tier classification table and develop a model selection handbook to assist users in selecting the most cost-effective models for their specific requirements (Section 3.7).
 3.1 Overall Performance
@@
 Findings Consistent with Common Sense:
-(1) Undoubtedly, DeepSeek-V3 and DeepSeek-R1 demonstrate the best overall performance among all evaluated models..
-(2) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts. After 4-bit quantization, the model’s scores decrease by an average of 3.52%.
-(3) Within the same series, models follow the scaling law [5, 9] regardless of whether they have undergone reasoning enhancement.
-(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 178.74% and 54.36%, respectively.
+(1) Undoubtedly, DeepSeek-V3 and DeepSeek-R1 demonstrate the best overall performance among all evaluated models.
+The 4-bit quantized version of DeepSeek-R1 demonstrates only a 0.84% performance degradation compared to the full-precision version, while still outperforming both Qwen and Llama series models.
+(2) Overall, reasoning-enhanced models outperform their original instruction-tuned counterparts. After 4-bit quantization, the model’s performance decreases by an average of 1.93%.
+(3) Overall, given identical model architectures and training data, larger parameter models demonstrate superior performance, aligning with the scaling law [5, 9].
+(4) Distillation brings the most significant improvements to Qwen2.5-Math-1.5B and Qwen2.5-Math-7B, with score increases of 212.47% and 68.19%, respectively.
 This is because they are math-focused and perform poorly on general tasks, but the distilled reasoning data from DeepSeek-R1 significantly enhances their general capabilities.
@@
 The average scores of models on each task.
-The "Instruct Model" refers to the original Instruction-tuned models without reasoning enhancement, the "Reasoning Enhancement Model" refers to reasoning-enhanced models that have been enhanced using DeepSeek-R1’s reasoning data, and the "Quantization Model" refers to 4-bit quantized version of the reasoning enhancement models.
+The "Instruct Model" refers to the original Instruction-tuned models without reasoning enhancement, the "Reasoning Enhancement Model" refers to reasoning-enhanced models that have been distilled using DeepSeek-R1’s reasoning data or the native reasoning models, and the "Quantization Model" refers to 4-bit quantized version of the reasoning enhancement models.
 Findings Contrary to Common Sense:
@@
 The counterintuitive result is likely due to the predominance of Chinese data in our evaluation dataset, and this phenomenon also appears later in the discussion of performance comparison by task (Section 3.2) and by subtask (Section 3.3).
-(6) Qwen-2.5-14B exhibits performance degradation after distillation.
+(6) Smaller models can achieve enhanced capabilities when employing optimized training strategies and higher-quality data.
+For instance, the reasoning model QwQ-32B outperforms the original instruction-tuned model Qwen2.5-32B-Instruct by 3.62% in overall performance, surpasses the distilled model DeepSeek-R1-Distill-Qwen-32B by 2.56%, and trails the DeepSeek-R1 by only 1.57%.
 3.2 Performance Comparison By Task
 Figure 1(b) to  Figure 1 (f) compare the scores of the models on the five major task categories.
-Findings Consistent with Common Sense:
+Findings:
 (7) The two mathematical models achieve comparable performance in Logical Reasoning and Task Planning but under-perform in other tasks.
-(8) For Logical Reasoning tasks, all models show improvement after distillation, with the increase being greater than for other tasks.
-Additionally, the performance degradation of 4-bit quantization on Logical Reasoning tasks is the most significant, at 7.91%.
+(8) Compared to instruction-tuned model DeepSeek-V3, reasoning model DeepSeek-R1 achieves significant gains of 5.4% and 3.0% in logical reasoning and task planning tasks, respectively.
+However, it demonstrates performance decreases of 2.1% and 1.6% in text understanding and generation tasks.
+(9) When distilled using DeepSeek-R1’s data, Qwen and Llama model families show overall performance improvements of 1.63% and 2.65% respectively.
+Particularly in logical reasoning tasks, the distilled models achieve substantial gains of 9.50% and 19.92% respectively.
+However, Performance degradation occurs after distillation in: Text Understanding (Llama-3.1-8B, Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Llama-3.1-8B, Qwen2.5-14B, DeepSeek-V3), and Task Planning (Llama-3.3-70B).
+(10) Quantization affects different task types variably: Minimal impact on text generation and task planning (average performance drop: 0.32%); Moderate impact on text comprehension and information extraction (average drops: 1.58% and 1.09%); Significant impact on logical reasoning (average drop: 6.49%)
 Figure 2: Average scores of the DeepSeek series models on the 27 subcategories.
-Findings Contrary to Common Sense:
-(9) Performance degradation occurs after distillation in: Text Understanding (Qwen2.5-14B, Qwen2.5-32B, and DeepSeek-V3), Information Extraction (Llama-3.1-8B, Qwen2.5-14B), and Text Generation (Qwen2.5-14B, DeepSeek-V3), and Task Planning (Llama-3.1-8B, Llama-3.3-70B).
-(10) For Logical Reasoning tasks, original Llama-3.3-70B underperforms the smaller model Qwen2.5-14B.
-After distillation, DeepSeek-R1-Distilled-Llama-3.3-70B surpasses DeepSeek-R1-Distilled-Qwen2.5-32B.
-(11) For Information Extraction tasks, Qwen2.5-32B consistently outperform the larger model Llama-3.3-70B both before and after distillation.
 3.3 Performance Comparison By Subtask
@@
 After analyzing these results, here are some interesting findings:
-(12) DeepSeek models dominate 23/27 subtasks, except for Short Text Classification, Named Entity Recognition, and Common Sense Question Answering.
-(13) DeepSeek-R1 shows relative weaknesses compared to DeepSeek-V3 in Long Text Classification (-11.3%), Part-of-Speech Tagging (-13.4%), and Open-source Question Answering (-17.05%).
-(14) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 31.45%.
+(11) DeepSeek models achieve dominant performance in 21 out of 27 subtasks.
+However, QwQ-32B demonstrates superior capability in the remaining six subtasks: Named Entity Recognition, Event Extraction, Common Sense QA, Open-domain QA, Semantic Entailment Recognition, and Code Generation.
+(12) Compared to other tasks, distillation brings the highest gains in Complex Mathematical Computation subtask, with an average improvement of 62.86%.
 3.4 Performance Comparison By Model
@@
 Figure 3: Performance of each model group on five major tasks.
-3.5 Model Selection Guidance for Users
-To assist users in selecting the most suitable DeepSeek model based on their specific application requirements, we present the evaluation results using line charts, following the A-Eval methodology.
+3.5 Line Chart Results
+we present the evaluation results using line charts, following the A-Eval methodology.
 These visualizations provide a clear and intuitive comparison of model performance across different tasks and subcategories.
 Figure 4: Line chart for model selection on five major tasks.
-Figure 4 illustrates the scores of the 21 evaluated models on each major task, as well as their average scores across all tasks.
+Figure 4 illustrates the scores of the 22 evaluated models on each major task, as well as their average scores across all tasks.
 This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
-In Figure 5, we further break down the performance of the 21 models by showing their average scores within each of the five major task categories and their corresponding subcategories.
+In Figure 5, we further break down the performance of the 22 models by showing their average scores within each of the five major task categories and their corresponding subcategories.
 These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
 Figure 5: Line chart for model selection on five major tasks and corresponding subtasks.
-Using these capability boundary quantization curves, users can make informed decisions when selecting the most suitable DeepSeek model for their specific tasks and performance expectations.
+3.6 Performance Comparison By Difficulty Level
+In A-Eval and A-Eval-2.0, each QA pair is assigned a difficulty level label: Easy, Medium, or Hard.
+It is evident that models tend to achieve lower scores as the difficulty of the questions increases.
+However, there has been no quantitative analysis of how DeepSeek models perform on questions of varying difficulty levels.
+To address this gap, we separately analyze the evaluation results for each difficulty level and present our findings.
+Figure 6 (a) illustrates the average scores of 22 models across task categories for Easy, Medium, and Hard data.
+Meanwhile, Figure 6 (b) - (f) depict the scores of the 22 models for each task category across the three difficulty levels.
+Figure 6: Average performance of 22 models on Easy, Medium, and Hard data.
+In Figure 7, we analyze the percentage score gains brought by reasoning enhancements on Easy, Medium, and Hard data. An interesting observation emerges:
+(13) As the data difficulty increases, distillation or reasoning enhancements yield higher performance gains for the models.
+Interestingly, reasoning enhancements can even have a negative impact on simpler problems.
+Figure 7: Performance gain of reasoning enhancement on Easy, Medium, and Hard data.
+(a) Performance gain of each model group across all tasks.
+(b) Performance gain of each task category across all models (excluding math models).
+3.7 Model Selection Guidance for Users
+To assist users in selecting the most suitable model based on their specific application requirements, we classify model performance into five tiers (A+, A, B, C, D) based on their scores in the five major task categories, resulting in the capability boundary quantification table (Table 2).
 Table 2: Performance tier classification of DeepSeek models across major task categories.
-Models are categorized into tiers (A+: >85, A: 80-85, B: 75-80, C: 65-75, D: <65) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
+Models are categorized into tiers (A+: >85, A: 80-85, B: 75-80, C: 60-75, D: <60) based on their scores in Text Understanding (TU), Information Extraction (IE), Text Generation (TG), Logical Reasoning (LR), and Task Planning (TP).
 Model Model Scale TU IE TG LR TP
@@
 Qwen2.5-Math-7B-Instruct 7B D D D D D
-DeepSeek-R1-Distill-Qwen-7B 7B C D D B C
+DeepSeek-R1-Distill-Qwen-7B 7B D D D C C
 Distill-Qwen-7B-Q4KM 7B D D D C C
-Llama-3.1-8B-Instruct 8B C D C D B
-DeepSeek-R1-Distill-Llama-8B 8B C D C C C
-Distill-Llama-8B-Q4KM 8B C D C D C
-Qwen2.5-14B-Instruct 14B B B A B C
-DeepSeek-R1-Distill-Qwen-14B 14B B C B A B
-Distill-Qwen-14B-Q4KM 14B B C A C B
+Llama-3.1-8B-Instruct 8B C C C D C
+DeepSeek-R1-Distill-Llama-8B 8B C C C C C
+Distill-Llama-8B-Q4KM 8B D C C D C
+Qwen2.5-14B-Instruct 14B B B A C C
+DeepSeek-R1-Distill-Qwen-14B 14B C C A A B
+Distill-Qwen-14B-Q4KM 14B C C A C B
 Qwen2.5-32B-Instruct 32B B B A B C
-DeepSeek-R1-Distill-Qwen-32B 32B B B A B A
-Distill-Qwen-32B-Q4KM 32B B B A+ B B
-Llama-3.3-70B-Instruct 70B C C C C B
-DeepSeek-R1-Distill-Llama-70B 70B B C B A C
-Distill-Llama-70B-Q4KM 70B B C B C C
-DeepSeek-V3 671B|37B A B A+ A A+
+DeepSeek-R1-Distill-Qwen-32B 32B B B A A B
+Distill-Qwen-32B-Q4KM 32B B B A A B
+Llama-3.3-70B-Instruct 70B A B A B B
+DeepSeek-R1-Distill-Llama-70B 70B A A A A A
+Distill-Llama-70B-Q4KM 70B B A B A A
+DeepSeek-V3 671B|37B A A A+ A A+
 DeepSeek-R1 671B|37B A A A A+ A+
-DeepSeek-R1-Q4KM 671B|37B B C A A+ A+
-To further simplify the model selection process for users without extensive knowledge of LLMs, we classify the performance of DeepSeek models into five tiers (A+, A, B, C, D) based on their scores in the five major task categories.
+DeepSeek-R1-Q4KM 671B|37B A A A A A+
+QwQ-32B 32B A A A A A
 As shown in Table 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65).
-This tiered classification provides an intuitive and accessible way for users to quickly identify the most suitable models for their specific application needs, balancing performance and cost-effectiveness.
-Based on Table 2, we provide specific examples of model selection to illustrate its practical utility:
-If users require A+ level performance, the following models are recommended for each task category ((highlighted in blue in Table 2)):
-Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
-Information Extraction: DeepSeek-R1 (A)
-Text Generation: Distill-Qwen-32B-Q4KM (A+)
-Logical Reasoning: DeepSeek-R1-Q4KM (A+)
-Task Planning: DeepSeek-R1-Q4KM (A+)
-If users require A level performance, the following models are recommended for each task category (highlighted in red in Table 2):
-Text Understanding: DeepSeek-V3 (A) or DeepSeek-R1 (A)
-Information Extraction: DeepSeek-R1 (A)
-Text Generation: Distill-Qwen-14B-Q4KM (A)
-Logical Reasoning: DeepSeek-R1-Distill-Qwen-14B (A)
-Task Planning: DeepSeek-R1-Distill-Qwen-32B (A)
-For users with more flexible requirements, such as B level performance, the following models offer a cost-effective balance (highlighted in green in Table 2):
-Text Understanding: Distill-Qwen-14B-Q4KM (B)
-Information Extraction: Qwen2.5-14B-Instruct (B)
-Text Generation: Distill-Qwen-14B-Q4KM (A)
-Logical Reasoning: DeepSeek-R1-Distill-Qwen-7B (B)
-Task Planning: Distill-Qwen-14B-Q4KM (B)
-These examples demonstrate how the tiered classification can guide users in selecting models that meet their specific performance and cost requirements, ensuring optimal deployment in real-world applications.
-3.6 Performance Comparison By Difficulty Level
-In A-Eval and A-Eval-2.0, each QA pair is assigned a difficulty level label: Easy, Medium, or Hard.
-It is evident that models tend to achieve lower scores as the difficulty of the questions increases.
-However, there has been no quantitative analysis of how DeepSeek models perform on questions of varying difficulty levels.
-To address this gap, we separately analyze the evaluation results for each difficulty level and present our findings.
-Figure 6 (a) illustrates the average scores of 21 models across task categories for Easy, Medium, and Hard data.
-Meanwhile, Figure 6 (b) - (f) depict the scores of the 21 models for each task category across the three difficulty levels.
-Figure 6: Average performance of 21 models on Easy, Medium, and Hard data.
-In Figure 7, we present the scores of 21 models on Easy, Medium, and Hard data for each sub-task category.
-Figure 7: Performance of 21 models on Easy, Medium, and Hard data for each task and sub-task.
-In Figure 8, we analyze the percentage score gains brought by reasoning enhancements on Easy, Medium, and Hard data. An interesting observation emerges:
-(14) As the difficulty of the data increases, distillation or reasoning enhancements yield higher performance gains for the models.
-Interestingly, reasoning enhancements can even have a negative impact on simpler problems.
-Figure 8: Performance gain of reasoning enhancement on Easy, Medium, and Hard data.
-(a) Performance gain of each model group across all tasks.
-(b) Performance gain of each task category across all models (excluding math models).
-It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria.
+This tiered classification provides an intuitive and accessible way for users to quickly identify models’ performance for specific abilities.
+To further simplify the model selection process for users without extensive knowledge of LLMs, we comb comprehensive real-world applications and develop the model selection handbook that is shown in Figure 8.
+Figure 8: The model selection handbook.
+The model selection handbook illustrates the relation among models, their capabilities, and practical applications, enabling users to select the most cost-effective model without any effort.
+It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, the characteristics of data distribution, and the setting of evaluation criteria.
 Therefore, our evaluation results may not be the most objective, comprehensive, or accurate.
-As we will continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
+As we continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
 4 Conclusion
-To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models, their distilled variants, and 4-bit quantized models on the modified A-Eval-2.0 benchmark.
+To better understand how DeepSeek models perform in real-world applications, we comprehensively evaluate DeepSeek models, their distilled variants, 4-bit quantized models, and QwQ-32B on the modified A-Eval-2.0 benchmark.
 Our analysis reveals that reasoning-enhanced models, while generally powerful, are not universally superior across all tasks.
-Finally, we quantify the capability boundary of DeepSeek models through performance tier classifications and intuitive line charts.
-Specific examples provide actionable insights to help users select and deploy the most cost-effective DeepSeek models based on their specific application requirements.
+Additional, we quantify the capability boundary of DeepSeek models through performance tier classifications.
+Finally, we develop the model selection handbook, providing actionable insights to help users select and deploy the most cost-effective models based on their specific application requirements.
 In the future, we will continue to advance this work by optimizing evaluation benchmarks and promptly updating evaluation results to the community.

### v5

@@
-Corresponding author(s)
+
 Abstract
@@
 However, detailed evaluations for DeepSeek Series models from the perspective of real-world applications are lacking, making it challenging for users to select the most suitable DeepSeek models for their specific needs.
-To address this gap, we conduct a systematic evaluation of the DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, their corresponding 4-bit quantized models, and the reasoning model QwQ-32B using the enhanced A-Eval benchmark, A-Eval-2.0.
-Through a comparative analysis of original instruction-tuned models and their distilled counterparts, we investigate how reasoning enhancements impact performance across diverse practical tasks.
-Our results show that reasoning-enhanced models, while generally powerful, do not universally outperform across all tasks, with performance gains varying significantly across tasks and models.
-To assist users in model selection, we quantify the capability boundary of DeepSeek models through performance tier classifications.
-Based on the quantification results, we develop a model selection handbook that clearly illustrates the relation among models, their capabilities and practical applications.
-This handbook enables users to select the most cost-effective models without efforts, ensuring optimal performance and resource efficiency in real-world applications.
+To address this gap, we presents the first comprehensive evaluation of the DeepSeek and its related models (including DeepSeek-V3, DeepSeek-R1, DeepSeek-R1-Distill-Qwen series, DeepSeek-R1-Distill-Llama series, their corresponding 4-bit quantized models, and the reasoning model QwQ-32B) using our enhanced A-Eval benchmark, A-Eval-2.0.
+Our systematic analysis reveals several key insights: (1) Given identical model architectures and training data, larger parameter models demonstrate superior performance, aligning with the scaling law.
+However, smaller models may achieve enhanced capabilities when employing optimized training strategies and higher-quality data; (2) Reasoning-enhanced model show significant performance gains in logical reasoning tasks but may underperform in text understanding and generation tasks; (3) As the data difficulty increases, distillation or reasoning enhancements yield higher performance gains for the models.
+Interestingly, reasoning enhancements can even have a negative impact on simpler problems; (4) Quantization impacts different capabilities unevenly, with significant drop on logical reasoning and minimal impact on text generation.
+Based on these results and findings, we design an model selection handbook enabling users to select the most cost-effective models without efforts.
 It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results.
@@
 Please refer to the latest version of the paper for the most current results and conclusions.
-Keywords:
-DeepSeek application-driven evaluation reasoning enhancements capability boundary model selection handbook
 1 Introduction
@@
 Third, is the extent of performance loss observed in the quantized model when applied to various real-world tasks within an acceptable range?
-To address the gap, we first create A-Eval-2.0 by modifying some QAs in A-Eval [10], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories.
+To address the gap, we first create A-Eval-2.0 by modifying some QAs in our previous A-Eval [10], an application-driven evaluation benchmark comprising 678 human-curated question-answer (QA) pairs across five major task categories (Text Understanding, Information Extraction, Text Generation, Logical Reasoning, and Task Planning) and 27 subcategories.
 Then, we evaluate DeepSeek models on A-Eval-2.0.
@@
 These results and findings help guide users in selecting the most suitable DeepSeek models for their specific needs, enabling cost-effective deployment in real-world applications.
+Based on these results and findings, we contribute: (a) The open-sourced A-Eval-2.0 benchmark, including dataset and an evaluation toolkit; (b) The capability boundary quantification results for DeepSeek models; (c) An model selection handbook enabling users to select the most cost-effective models without efforts.
+Our work provides both methodological contributions to LLM evaluation and practical guidance for real-world applications.
 It is important to emphasize that the selection of test samples and the design of evaluation criteria will inevitably introduce certain biases into the evaluation results.
@@
 2.2 Dataset
-The A-Eval [10] benchmark, designed for practical application scenarios, quantifies the capability boundaries of LLMs through 678 manually curated QA pairs across three difficulty level, five major categories and 27 subcategories.
-To accommodate the enhanced reasoning capabilities of the DeepSeek series models, we refine the sub-tasks and QAs of Task Planning category in A-Eval, thereby constructing the evaluation dataset A-Eval-2.0 utilized in this evaluation.
+The A-Eval benchmark [10] was originally designed to assess the capability boundaries of large language models (LLMs) in practical application scenarios.
+It consists of 678 manually curated question-answer pairs spanning three difficulty levels, five major categories, and 27 subcategories.
+While comprehensive for its initial purpose, A-Eval was primarily designed for evaluating conventional instruction-tuned models across different parameter scales.
+However, the original benchmark has several limitations in the context of modern LLM development:
+1. It does not account for specialized model variants like distilled, quantized, or reasoning-optimized architectures.
+2. Its evaluation granularity is insufficient for complex reasoning and planning tasks.
+3. The difficulty progression of data lacks precision in assessing advanced cognitive capabilities.
+To address these limitations and better evaluate the enhanced reasoning capacities of DeepSeek models, we introduce A-Eval-2.0.
+This enhanced version maintains the core structure of A-Eval while incorporating critical improvements in two key aspects:
+Extended Task Planning Subcategories: We add "Scheme Optimization" as a new subcategory under Task Planning category and make 40 new test QA pairs, focusing on evaluating models’ ability to execute multi-step plans under constraints and optimize the plans according to dynamic feedback.
+This sub-task can more fully test the complex planning ability of the reasoning models.
+Refined Logical Reasoning and Task Planning Categories: We refine Logical Reasoning and Task Planning ability into Primary, Intermediate and Advanced levels and make clear definitions:
+Primary Logical Reasoning: Processes well-defined logical verification or contradiction detection tasks, typically accomplished through simple rules and single-step causal inference.
+Intermediate Logical Reasoning: Accomplishes multi-factor, multi-step causal reasoning within specific domains by processing conventional constraints and establishing localized logical chains.
+Advanced Logical Reasoning: Addresses domain-specific complex reasoning challenges requiring professional knowledge integration for comprehensive systemic deduction, backtracking, and refinement, ultimately generating professional-grade decisions or solutions.
+Primary Task Planning: Formulates linearly decomposable deterministic task flows through standardized breakdown of finite steps, executing basic commonsense planning with either explicit parallel branches or no parallelism, without requiring dynamic adjustment.
+Intermediate Task Planning: Designs multi-constrained comprehensive solutions for single entities by balancing resources, priorities, and personalized needs, while handling common limited variants and optimizing execution paths.
+Advanced Task Planning: Constructs general-purpose intelligent agents capable of professional-grade, system-level planning in dynamic multi-agent environments, integrating real-time environmental variables to adapt execution strategies.
 We use A-Eval-2.0’s dataset as the evaluation dataset to analyze the performance of DeepSeek series models in practical applications.
@@
 2.3 Evaluation Process
-We follow the automatic and manual scoring methods of A-Eval, and the automatic scoring process includes:
+We follow the automatic scoring pipeline of original A-Eval, and the automatic scoring process includes:
 Inference: Feed each question ${{Q_{i}}}$ into the evaluated model to generate a prediction ${{P_{i}}}$.
@@
 Scoring: Combine the prompt and triplet$\left({{Q_{i}},{A_{i}},{P_{i}}}\right)$, input them into the scoring model, and obtain the scoring output ${{S_{i}}}$ between 0 and 100.
-We use A-Eval’s scoring prompt and a greater LLM, Qwen2.5-72B-Instruct as the scoring model.
+To better evaluate the enhanced reasoning capacities of DeepSeek models, we optimize the scoring details in several aspects:
+Advanced Scoring Model: We employ DeepSeek-R1 as the scoring model, which demonstrates superior judgment capabilities compared to original instruction-tuned model.
+Optimized Scoring Prompt: To better assess the capabilities of instruction-tuned models, distilled models, reasoning models and quantized models, we refine the scoring prompt.
+Two-Phase Expert Review: To ensure the reliability of the evaluation results, our evaluation expert team conducts a two-round review of the automatic scoring results:
+Phase 1.
+Independent Reviewing: Firstly, $N$ evaluation experts independently assess each automated scoring tuple $\left({{Q_{i}},{A_{i}},{P_{i}},{S_{i}}}\right)$ after aligning expert scoring standards.
+Each expert may output the manual score $E_{i}^{k}={S_{i}}$ if the automated score ${{S_{i}}}$ is deemed reasonable or provide the manual score $E_{i}^{k}\neq{S_{i}}$, where $k=\left\{{1,2,....,N}\right\}$.
+For each tuple, we then calculate the consensus ratio: $C{R_{i}}={1\over N}\sum\limits_{k=1}^{N}{{\rm I}\left({E_{i}^{k}={S_{i}}}\right)}$ where ${\rm I}$ denotes the indicator function.
+Finally, ${S_{i}}$ is considered reliable if $C{R_{i}}\geq\tau$ (we set $N=5$ and $\tau=0.8$), indicating at least 80% expert agreement with the automated score ${S_{i}}$.
+Phase 2.
+Consensus Reconciliation: For tuples failing to meet the consensus threshold ($C{R_{i}}<\tau$), the evaluation team conducts a focused group discussion.
+Following discussion, each expert submits a revised score $E_{i}^{k*}$.
+The final revised score $S_{i}^{*}$ is computed using averaging method that discards the highest and lowest scores.
+This two-phase review process ensures both independent assessment and collective wisdom are incorporated into the final evaluation results. The optimized scoring prompt is as follows:
+Scoring Prompt Template
 3 Results and Discussion
@@
 Then, we statistic performance of models on datasets with three difficulty levels—easy, medium, and hard—and discuss models’ performance to variations in question difficulty (Section 3.6).
-Finally, we quantify models’ capability boundary through a tier classification table and develop a model selection handbook to assist users in selecting the most cost-effective models for their specific requirements (Section 3.7).
+Then, we quantify models’ capability boundary through a tier classification table, illustrating performance levels of models on different capabilities (Section 3.7).
+Finally, develop a model selection handbook to assist users in selecting the most cost-effective models for their specific requirements without efforts (Section 3.8).
 3.1 Overall Performance
@@
 These visualizations provide a clear and intuitive comparison of model performance across different tasks and subcategories.
-Figure 4: Line chart for model selection on five major tasks.
-Figure 4 illustrates the scores of the 22 evaluated models on each major task, as well as their average scores across all tasks.
+Figure 4: Line charts for evaluation performance. (a) Line chart on five major tasks. (b) - (f) Line charts on five major tasks and corresponding subtasks
+Figure 4 (a) illustrates the scores of the 22 evaluated models on each major task, as well as their average scores across all tasks.
 This comprehensive overview allows users to quickly identify models that excel in specific major task types or demonstrate strong overall performance.
-In Figure 5, we further break down the performance of the 22 models by showing their average scores within each of the five major task categories and their corresponding subcategories.
+In Figure 4 (b) to Figure 4 (f), we further break down the performance of the 22 models by showing their average scores within each of the five major task categories and their corresponding subcategories.
 These detailed visualizations enable users to assess model capabilities at a granular level, ensuring a precise match between model strengths and task requirements.
-Figure 5: Line chart for model selection on five major tasks and corresponding subtasks.
 3.6 Performance Comparison By Difficulty Level
@@
 Figure 5: Average performance of 22 models on Easy, Medium, and Hard data.
-In Figure 7, we analyze the percentage score gains brought by reasoning enhancements on Easy, Medium, and Hard data. An interesting observation emerges:
-(13) As the data difficulty increases, distillation or reasoning enhancements yield higher performance gains for the models.
-Interestingly, reasoning enhancements can even have a negative impact on simpler problems.
-Figure 7: Performance gain of reasoning enhancement on Easy, Medium, and Hard data.
-(a) Performance gain of each model group across all tasks.
-(b) Performance gain of each task category across all models (excluding math models).
-3.7 Model Selection Guidance for Users
-To assist users in selecting the most suitable model based on their specific application requirements, we classify model performance into five tiers (A+, A, B, C, D) based on their scores in the five major task categories, resulting in the capability boundary quantification table (Table 2).
 Table 2: Performance tier classification of DeepSeek models across major task categories.
@@
 QwQ-32B 32B A A A A A
+In Figure 6, we analyze the percentage score gains brought by reasoning enhancements on Easy, Medium, and Hard data. An interesting observation emerges:
+(13) As the data difficulty increases, distillation or reasoning enhancements yield higher performance gains for the models.
+Interestingly, reasoning enhancements can even have a negative impact on simpler problems.
+Figure 6: Performance gain of reasoning enhancement on Easy, Medium, and Hard data.
+(a) Performance gain of each model group across all tasks.
+(b) Performance gain of each task category across all models (excluding math models).
+3.7 Model Performance Tier Classifications
+We classify model performance into five tiers (A+, A, B, C, D) based on their scores in the five major task categories, resulting in the capability boundary quantification table (Table 2).
 As shown in Table 2, models are assigned a tier for each task category, with A+ representing the highest performance (scores > 85) and D indicating the lowest (scores < 65).
 This tiered classification provides an intuitive and accessible way for users to quickly identify models’ performance for specific abilities.
-To further simplify the model selection process for users without extensive knowledge of LLMs, we comb comprehensive real-world applications and develop the model selection handbook that is shown in Figure 8.
-Figure 8: The model selection handbook.
-The model selection handbook illustrates the relation among models, their capabilities, and practical applications, enabling users to select the most cost-effective model without any effort.
-It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, the characteristics of data distribution, and the setting of evaluation criteria.
-Therefore, our evaluation results may not be the most objective, comprehensive, or accurate.
-As we continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
+3.8 Model Selection Notebook for Users
+To assist users in selecting the most suitable model based on their specific application requirements, we develop a model selection notebook consisting of four components.
+1.
+LLM Capability Definitions and Application Scenarios.
+As shown in Figure 7, we systematically categorize a substantial number of real-world application scenarios and analyze the specific model capabilities required for each.
+Figure 7: LLM capability definitions and application scenarios, including 15 well-defined model capabilities along with their corresponding 105 application scenarios.
+2.
+LLM “Parameter Scale-Capability-Application Scenarios” Relation.
+As shown in Figure 8, we analyze the relationship among parameter scale, capability, and application based on the evaluation results and the model performance tier classifications in Table 2.
+Figure 8: LLM “Parameter Scale-Capability-Application” relation.
+The figure illustrates the correlations among models, their capabilities, and practical applications.
+For each capability, the most cost-effective model is indicated by green lines, while other models with higher computational costs are shown in red.
+3. Supported Application Scenarios for Each Model. As shown in Figure 9, we clearly demonstrate the specific application scenarios that each model can successfully accomplish.
+Figure 9: Supported application scenarios of each model.
+4.
+Cost-Effective Model Recommendations for Target Application Scenarios.
+As illustrated in Figure 10, to simplify the model selection process for users without extensive knowledge of LLMs, we provide cost-effective model recommendations for each application scenario, along with detailed hardware deployment requirements.
+Figure 10: Cost-Effective model recommendations for target application scenarios.
+Using our model selection notebook, users can efficiently identify the most cost-effective models, ensuring optimal performance and resource efficiency in real-world applications.
+It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results.
+We will continuously optimize the evaluation benchmarks and periodically update this paper to provide more comprehensive and accurate evaluation results.
+Please refer to the latest version of the paper for the most current results and conclusions.
 4 Conclusion